Liza is four times as old as Nene. Six years ago, she was 7 times as old as Nene. Find their present age.
Answer by mananth(16946)

Nene x years
Liza 4x years
...
6 years ago
Nene = x-6
Liza = 4x-6
...
4x-6=7(x-6)
4x-6=7x-42
7x-4x=42-6
3x=36
/3
x=12 Nene's age
..
Liza = 4x = 48 years